OpenStudy (anonymous):

please explain how you get your answer 3x+4y=7 x=3-2y

OpenStudy (anonymous):

Where did you get the -14y from?

OpenStudy (anonymous):

Oh i see

OpenStudy (ybarrap):

3x+4y=7 x=3-2y 4y = 7 - 3x (*) 2y = 3 - x (**) (*) - 3 time (**) 4y - 3(2y) = 7 - 3(3) - 3x - 3(-x) 4y - 6y = -2 -2y = -2 y = -2/-2 = 1 3x = 7 - 4y = 7 - 4(1) = 3 x = 1 Ans x = 1, y = 1

OpenStudy (anonymous):

sorry made a calculation mistake up there.

OpenStudy (anonymous):

Why did you change the problem around?